OpenStudy (anonymous):

Write the equation y = -x2 – 3x in vertex form. What is the vertex and the axis of symmetry?

OpenStudy (anonymous):

y = -x2 – 3x y = -x2 – 2*(3/2) x -9/4 +9/4 (y-9/4)=-(x-3/2)^2 vertex 3/2,9/4 and axis is -ve direction of x=3/2

OpenStudy (anonymous):

thanx again